OpenStudy (anonymous):

a to the power of 1/4 and b to the power of 6 multiplied by a to the power of 3 and b to the power of 1/4 divided by a to the power of 1/2 and b to the power of 1/6

OpenStudy (campbell_st):

so the problem is \[\frac{a^{\frac{1}{4}}b^6 \times a^3b^{\frac{1}{4}}}{a^{\frac{1}{2}}b^{\frac{1}{6}}}\] so the rules you need when multiplying the same base add the powers when dividing, subtract the powers so you will get \[\frac{a^{\frac{1}{4} + 3}b^{6 + \frac{1}{4}}}{a^{\frac{1}{2}}b^{\frac{1}{6}}}\] so you get \[\frac{a^{\frac{13}{4}}b^{\frac{25}{4}}}{a^{\frac{1}{2}}b^{\frac{1}{6}}}\] now do the division and subtract the powers hope it helps

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

do you subtract the a's or b's?